Mocha company manufactures a single product by a continous process, involving three production departments. The records indicate that direct materials, direct labor, and applied factory overhead for department 1 were $100,000, $125,000, and $150,000, respectively. The records further inicate that direct materials, direct labor, and applied given the following cost and activity observations for bounty company's utilities, use the high-low method to calculate bounty's variable utilities cost per machine our. Rounds to the nearest cent. cost machine hours March 3,100 15,000 April 2,700 10,000 May 2,900 12,000 June 3,600 18,000 answer A. $10.00 B. $.0.63 C. $0.14 D. $0.67
Answer #1

Variable cost per machine hour = (3600-2700)/(18000-10000)=900/8000=0.1125

3600-highest cost of activity

2700-lowest cost of activity

18000-highest level of activity

10000-lowest level of activiity

comment for any clarification